Fermi National Laboratory


Accelerator Update for December 5, 2005 - December 7, 2005

Monday December 5

The day shift began with the Tevatron (TeV) sector B1 separator bake out continuing, and with MiniBooNE, NuMI, and the Meson MCenter experiments taking beam.

Experts accessed the Antiproton Source (Pbar) Debuncher area. The TeV B1 cryo system cooldown began early in the afternoon. The Main Injector (MI) conducted studies. Early in the evening, a power supply expert had to repair a Debuncher quadrupole (D:Q916). Later in the evening, Pbar came out of access and began reverse proton studies.

Tuesday December 6

The midnight shift began with TeV B1 cooldown and separator bake out continuing, with Pbar conducting reverse proton studies, and with MiniBooNE, NuMI, and MCenter taking beam.

Around 3 AM, the MC7 interlocks dropped. Operators re-interlocked the area and restored beam.

During the day, Pbar conducted reverse proton studies. NuMI hit a milestone with an integrated total of 1E20 protons on their target. A LCW leak was found in CUB. Repairing the leak would require shutting down the TeV LCW system. Cryo system technicians conducted wet engine maintenance around the TeV. A MI expert conducted slipstacking studies.

Operations reported at 11:51 PM that operators had completed interlocking the TeV.

Wednesday December 7

The midnight shift began with TeV B1 cooldown and separator bake out continuing, with Pbar studies, and with MiniBooNE, NuMI, and MCenter taking beam.

At 1:58 AM, Operations reported a helium leak on the TeV sector A4 cryo system. An expert came and repaired the leak.

Operations reported at 3:16 AM that the B1 cryo system was cold.

Operators successfully completed a TeV hipot at 6:58 AM.

At 7:45 AM, a Linac quadrupole power supply (L:Q31) developed a noise problem for the first of two times in twelve minutes. Operators were able to normalize it both times.

The Plans for Wednesday

The plans are to get permission and sign offs to power the TeV this morning and then start powering up some devices later today. The heating portion of the separator bake out will be completed today, but the bake out equipment must be removed before the TeV can conduct studies with beam. On Thursday, experts will access the TeV to check on the separators, pull the equipment, and do other miscellaneous work.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 12/7/2005   email Fermilab